OpenStudy (anonymous):

What decimal is equivalent to 4/3/12

OpenStudy (studybird123):

4.25 if its 4 3/12

OpenStudy (anonymous):

you have to stop writing the question like that it's confusing you did that yesterday .. :/

OpenStudy (anonymous):

lol how am I suppose to write it

OpenStudy (anonymous):

4+3/12

OpenStudy (anonymous):

it looks like a date lol

OpenStudy (anonymous):

Noob

OpenStudy (studybird123):

ok batman noob then o_o

OpenStudy (anonymous):

\[4 \frac{ 3 }{ 12 }\]

OpenStudy (anonymous):

HOW DID U DO THAT !

OpenStudy (anonymous):

like what some white board or something I dont know of

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

use the equation button type in the fraction

OpenStudy (anonymous):

\[4 \frac{ 3 }{ 12 } \implies 4+\frac{ 3 }{ 12 }\]

OpenStudy (anonymous):

But people can mistake 4 3/12 like as multiplication.

OpenStudy (anonymous):

yea ^ true

OpenStudy (anonymous):

\[4\frac{ 3 }{ ? }\]

OpenStudy (anonymous):

fail...

OpenStudy (anonymous):

But anyway 4.25 is correct

OpenStudy (anonymous):

lol k

OpenStudy (studybird123):

ok yay!